linda and dave leave simultaneously from the same starting point biking in opposite directions..linda bikes at 5 miles per hour and dave bikes at 8 miles per hour...how long will it be until they are 20 miles apart from each other?
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

moving apart at 13 mph
t=D/r
t=20/13
t=1 hr 32.5 min

Answer by checkley77(12844)

5X+8X=20
13X=20
X=20/13
X=1.538 HOURS THEY WILL BE 20 MILES APART.